OpenStudy (anonymous):

\[\sqrt[3]{270x}-\sqrt[3]{10x^4}\]

OpenStudy (anonymous):

\[\sqrt[3]{270x}-\sqrt[3]{10x^4}=\sqrt[3]{3^3*10x}-\sqrt[3]{10x*x^3}\] Simplified: \[3\sqrt[3]{10x}-x\sqrt[3]{10x}=\sqrt[3]{10x}(3-x)\]